OpenStudy (anonymous):

write a polynomial funtion of least degree with an intergal coefficants that has the given zeros, 1, 3, 0

myininaya (myininaya):

f(x)=(x-1)(x-3)(x-0)

OpenStudy (anonymous):

How did you get this please

myininaya (myininaya):

what makes the function zero?

myininaya (myininaya):

its zeros right?

OpenStudy (anonymous):

yes

myininaya (myininaya):

if i said the polynomial has zeros 2 and 3, then f(x)=(x-2)(x-3) because writing this will make 2 and 3 zeros of f.

OpenStudy (anonymous):

I LOVE YOU thank you

myininaya (myininaya):

if i said the polynomial has zeros -2,3,4,5, then f(x)=(x-(-2))(x-3)(x-4)(x-5)

myininaya (myininaya):

np